OpenStudy (anonymous):

A stack of bricks weighs 170 KN (Kilo-newtons). The stack exerts 200 KPa (Kilo-pascals) of pressure on the ground. What is the area upon which this pressure is exerted (in square ft)?

OpenStudy (anonymous):

P = F/A A= F/P A= 170x10^3N / 200 x 10^3Pa A= 0.85m^2 convert it into square foot..

OpenStudy (anonymous):

@bcmccants got it..??

OpenStudy (anonymous):

Yes, thank you. I was forgetting to square my numbers in the conversion.

OpenStudy (anonymous):

no problem.. :)